Shift Summary for Nov. 27 Day shift

Vivian O'Dell ((no email))
Sat, 27 Nov 1999 15:33:26 -0600

Shift Summary for Day shift, 11/27/99

On Shift: V. O'Dell, U. Erkarslan

Hours of beam delivered at all intensities: 8.5
Hours of beam delivered at nominal intensity: 8.5
Hours of 99 data (special runs not included): 7.75
Hours of special runs (laser scans, muon runs...): 0.25 (Pedestal Run)

Accounting of lost time:
"Boy, I'm full", tape swap.
Initializing runs.

Activities:
07:00 Day shift turns up to find owl shift struggling with
slow da. Day shift drinks coffee and waits patiently
for owl shift to fix all problems.

07:40 SDAQ complains DC4X_H_R/N out of limits. We page the
DC pager. There is no response so we page Leo. He
tells us not to worry about it. It is a known problem
and has been taken out of the trigger.

07:50 Angela calls sounding groggy. I think it took her a while
to realize that beeping noise was her pager. We told her
to go back to sleep.

07:50 While Leo has us on the phone he asks about SA4. SL looked
at the signals going into the FERA ADC's last night and
found that the SA4 pulse and the gate were slightly off,
which is causing a lower gain in the SA4. The veto seems
to be fine, though, so we keep on taking data.

10:55 Leo wanders in muttering about SA4. After futzing around
measuring gates and signals, he notices from online ped plots
that there is a 15 count ped shift in SA4.

11:00 We stop the run and Leo uses his persuasive powers to get
MCR to dump our beam so we can take a dedicated pedestal run.
They do/we do.

11:15 end of ped run.

11:30 start run 15068. SA4 distributions look much better.

12:35 onsplit requests mount UPM040. We do.

14:15 "Boy, I'm full". Taku ate too much turkey. We change
tapes.

14:30 Start run 15069.

15:30 Evening shift shows up.

=========================================================================
Runs ended this shift:
15066 BEAM_799 19M events
15067 pedestal run
15068 BEAM_799 5M events

Tapes brought to FCC this shift: 0

========================================================================
Follow up Notes on Previous Shift

gain drifts for 1476 and 2893 remain stable. More
channels are joining the drift hitlist. Does the
database need to be updated? We asked Ed about it when
he called in this morning, and he seemed confused but
said he would look into it.

========================================================================
Items to Follow up on Next Shift

keep an eye on the hcc inefficiencies.

========================================================================